Monthly spending ($)

 

 

 

 

 

 

Age

20s

30s

$0.00 - $49.99

75

29

$50.00 - $99.99

91

83

$100.00 - $149.99

113

90

$150.00 - $199.99

89

128

$200.00 - $249.99

63

159

$250.00 or more

34

46

Total customers

465

535

 

a. Given that a customer is in his/her 20s, what is the probability that he/she spends less than $100.00?



1
Expert's answer
2021-05-11T15:13:30-0400

The given data is-




(a) Total number of customers in 20s=465


Number of customers spends less than $100=75+91=166


Hence, Required Probability "=\dfrac{166}{465}=0.357"